OpenStudy (ttop0816):

is this right? The following system of equations is _____. y = -x + 2x + 6y = 4 independent equivalent <--- i chose this! inconsistent

OpenStudy (anonymous):

y = -x + 2x + 6y = 4

OpenStudy (anonymous):

first equation is incomplete ?

OpenStudy (ttop0816):

its -1/3x+2/3 !

OpenStudy (anonymous):

2x + 6y = 4 6y = -2x + 4 y = -x/3 + 2/3

OpenStudy (anonymous):

you're right